OpenStudy (anonymous):

The ratio of the legs of a right triangle is 5 to 12. If the hypotenuse is 39, what is the area?

OpenStudy (ash2326):

@Elak Do you know Pythagoras?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Let's assume one leg as x then other will be 12x/5 Do you agree?

OpenStudy (anonymous):

:)) ok

OpenStudy (anonymous):

but why?

OpenStudy (ash2326):

Let the other be y \[\frac x y= \frac 5 {12}\] Find y from this

OpenStudy (anonymous):

ok Y=5/12x :)

OpenStudy (ash2326):

Now use the pythagoras to find x, Can you do that?

OpenStudy (anonymous):

ok sorry , 12x=5y y=12x/5 :)

OpenStudy (ash2326):

oops :( I had missed that

OpenStudy (ash2326):

Could you find x?

OpenStudy (anonymous):

x^2+ (12x/5)^2= 39 isit right??? :((

OpenStudy (ash2326):

39^2

OpenStudy (ash2326):

@Elak Did you find x? it's easy

OpenStudy (anonymous):

x=15 ???

OpenStudy (ash2326):

Yes, I think you can find the area now?

OpenStudy (anonymous):

yes, yhank you very muuuch :)))

OpenStudy (ash2326):

you did everything :)

OpenStudy (anonymous):

but you help me, to think right ;)